OpenStudy (anonymous):

(7 + √2) (8 - 3√2) My answer is 50

OpenStudy (anonymous):

no this time you get middle terms

OpenStudy (anonymous):

I got 50 + 28√2

OpenStudy (anonymous):

\[(7+\sqrt{2})(8-3\sqrt{2})\] \[7\times 8-21\sqrt{2}+8\sqrt{2}-3\times 2\]

OpenStudy (anonymous):

the 50 is right, the middle ones add up to \(-13\sqrt{2}\)

OpenStudy (anonymous):

why are you multiplying by 8?

OpenStudy (anonymous):

because you wrote an 8 in the problem

OpenStudy (anonymous):

oh:) I understand now thanks!!!!! youre the beeeeeeeest

OpenStudy (anonymous):

yw (blush)

OpenStudy (anonymous):

:)))))!!!!!